...

Report from 8:30 meeting

by user

on
17

views

Report

Comments

Transcript

Report from 8:30 meeting
Thursday 15.7
● 04:20 Stable beams 13_8_8_8
–
–
–
–
–
–
Peak luminosity around 1.2E30 cm-2s-1.
Dampers on.
Intensity and luminosity lifetimes not as good as previous fill.
06:00 TOTEM RPs at 25 sigma vertical plane, 30 sigma hor. plane.
Problems with the TOTEM RP FESA server…
12:00 Damper off, lifetime of beam2 increases by a factor 2-3. Beam2 damper
settings not yet fully optimized.
– 13:00 Dump beam.
● 15:30 Injection – Hump hunting.
– Test effects of the GSM. No effect seen on the hump.
8:30 meeting
1
Lifetime damper ON - OFF
8:30 meeting
2
Hump vs. GSM signal in the tunnel
● Circulating beam (single nominal bunch) available at ~16:25
Started to switch OFF GSM network in the tunnel in the following order:
–
Sector 12 + TI2
Sector 23
Sector 34
sector 45
Sector 56
Sector 67
Sector 78
Sector 81 + TI8 + UX15
● By 17:24 all GSM transmitters in the tunnel were OFF
No significant modification on the hump amplitude
By 18:00 the GSM network was re-established (for a total interruption of 1.5 h)
This test was possible thanks to the contribution of IT/CS (A. Pascal, F. Chapron).
G. Arduini et al.
Hump vs. GSM signal in the tunnel
GSM ON – B2H
GSM OFF – B2H
GSM ON – B2V
GSM OFF – B2V
● No significant change in amplitude
● Typical slow drift in frequency as observed in other occasions.
Thursday 15.7
● 18:00 : Prepare for 13 + ½ nominal bunch configuration (LHCf)
– LBDS beam position interlocks reconfigured.
– Damper noise improvements for beam2.
– 20:00 Quench heater of MB in sector 45 discharging. To be fixed before the
week-end.
– 22:00 CMS reporting overpressure problem in their cavern.
● 23:00 Access for CMS
– LBDS beam position interlocks reverted to initial situation. Give up on 13 + ½
bunch scheme.
– QPS reset on S67. Not pre-cycled.
8:30 meeting
5
Friday 16.7
● 00:30 Injection
– Beams dumped twice by FMCM on RQ5.LR3. M. Zerlauth confirmed that the
problem seems to be internal to the FMCM + voltage divider system.
– Beam losses at injection, suspected scraping in the SPS, RF loops, energy
offsets…
– Abnormal bunch length increase in the LHC. Injection losses due to unbunched beam in the LHC. Large synchro loop errors.
● 04:00 RF work at injection.
– One explanation : the field in the LHC is off by around 0.3%, but not seen on
the trajectories. Still investigating…
● SPS : vacuum leak developing in sextant 4 since yesterday evening.
8:30 meeting
6
Bunch length and losses
8:30 meeting
7
Friday 16.7
●
●
●
●
Restore injection.
Multi-bunch injection.
Access for quench heater PS in S45.
Over the week-end:
– Physics
– Possibly short loss map measurements at injection.
8:30 meeting
8
Fly UP